OpenStudy (anonymous):

Write an appropriate direct variation equation if y = 27 when x = 9

hartnn (hartnn):

since its direct variation, u can write y=kx now put y=27,x=9 and find k....

OpenStudy (anonymous):

so 27=k(9) ?

hartnn (hartnn):

yup, so whats k ? k= 27/9 = ??

OpenStudy (anonymous):

3?

hartnn (hartnn):

yup, so the equation would be y=3x got it?

OpenStudy (anonymous):

yess! thhank you! :)

hartnn (hartnn):

welcome :)